### v2

@@ 4.1 Canonically decomposable matrices
As an example of canonical decomposability, for generic element $(A_{1},A_{2})$ in $(\real^{4\times 7})^{2}$, there exists a $(P,Q)\in GL_{4}\times GL_{7}$ such that $PA_{i}Q^{-1}=\oplus_{j=1}^{3}B_{ij}$, where $B_{i1},B_{i2}\in\real^{1\times 2}$ and $B_{i3}\in\real^{2\times 3}$ ([derksen2021], Example $2.8$). Here the term generic should be understood as an almost sure sense. An example with specific values of $A_{i}$’s and $(P,Q)$ is provided in Example $4$ of [maehara2011].
@@
The notion of canonical decomposability is mainly motivated by Kronecker quiver representation and its applications in the analysis of the sample size threshold for the existence of Kronecker MLE [derksen2021] (see [kac1980, kac1982] also). To say informally, the $n$-Kronecker quiver $Q$ is a directed acyclic graph of two vertices $x$ and $y$ with $n$ arrows. Then, a representation of $Q$ is to assign a finite-dimensional vector space to each vertex. If this representation cannot be written as a direct sum of a non-trivial subrepresentation, such a representation is referred to as a $\sigma$-stable representation. In the context of the Kronecker MLE problem, these vector spaces correspond to $\real{}^{p_{2}}$ and $\real{}^{p_{1}}$, and the arrows correspond to the $n$ data matrices. Using this Kronecker-quiver representation, along with the group-invariant theory, [derksen2021] characterized the scenarios of $(p_{1},p_{2},r)$ for which Kronecker MLE exists (see their Theorem $1.2$). For generic $(p_{1},p_{2})$, it turns out that $r>p_{1}/p_{2}+p_{2}/p_{1}$. Under this threshold, the uniqueness of Kronecker MLE also follows if it exists.
@@
To interpret this threshold, the canonical decomposability of $r$ data matrices in Definition 4.1 corresponds to whether the $r$ data matrices induce a $\sigma$-stable Kronecker quiver representation as arrows (see Section $3$ and $5$ of [derksen2021]). Then the threshold on $r$ for which Kronecker MLE exists comes from the minimum number of arrows for which the factors in $\real{}^{p_{1}}$ and $\real{}^{p_{2}}$ are well-connected. We shall mathematically formulate this below. We highlight that this decomposability notion also appears in other works on the sample size threshold analysis for Kronecker MLE. For example, [soloveychik2016] studied the threshold by analyzing the contribution of each block in the canonical decomposition of data matrices to the growth of the objective function $d$ in (2) (see Section $6.6$ of [soloveychik2016]). They referred canonically decomposable data matrices to as bad samples.
@@
In these works, the sample size threshold should be understood in a generic (almost sure) sense. If $r\geq p$ and $r$ data matrices are linearly independent, then the Kronecker MLE always uniquely exists [srivastava2008], not just generically. However, in a strict algebra sense, even if $r>p_{1}/p_{2}+p_{2}/p_{1}$ and the linear independence holds for data matrices, the Kronecker MLE may not exist as observed in Example 2.3. Nevertheless, note that the data matrices in that example are canonically decomposable. Thus, a natural question one could raise is whether any canonically decomposable matrices never admit the Kronecker MLE. It turns out that the answer is no, as shown in Example 4.2. This example also suggests that the canonically decomposable matrices are the singularities that may prevent the set of rank-$r$ cores from being a smooth manifold, in light of Sard’s theorem.
@@
Denote the subset of canonically decomposable matrices in $(\real^{p_{1}\times p_{2}})^{r}$ by $\mathcal{V}_{p_{1},p_{2},r}$. We show that this set is closed and has a Lebesgue measure zero. Note that its analogous results have been proven based on group-invariant theory and representation-theoretic approaches (see Proposition $3.19$ of [hoskins2012], and Lemma $2.16$ and Section $5$ of [derksen2021]). However, we provide a proof using a more direct language of algebraic geometry to make the article self-contained and better motivate the canonical decomposability in studying the fixed-rank core covariance manifold.
@@
Proof 4.4.
See Appendix A.3 for the proof.
Proposition 7.
@@
Proof 4.5.
See Appendix A.3 for the proof.
@@
Proof 4.7.
Suppose otherwise. Then there exists indecomopsable $A$ and $(P,Q)\in GL_{p_{1}}\times GL_{p_{2}}$ such that the graph $G_{A,P,Q}$ is disconnected. Hence, there exist partitions $U_{1}$ and $U_{2}$ of $\left\{s_{i}\right\}$ and accordingly $W_{1}$ and $W_{2}$ of $\left\{q_{j}\right\}$ such that a vertex in $U_{1}$ (resp. $U_{2}$) is never connected to $W_{2}$ (resp. $W_{1}$). After arranging the row and columns of $PA_{i}Q^{-1}$, we can obtain $(P^{\prime},Q^{\prime})\in GL_{p_{1}}\times GL_{p_{2}}$ such that $P^{\prime}A_{i}(Q^{\prime})^{-1}$ is of non-trivial block-diagonal form where the zero entries correspond to the absence of edges between $U_{1}$ (resp. $U_{2}$) and $W_{2}$ (resp. $W_{1}$), contradicting the indecomposability of $A$.
@@ 4.2 Proof of the smooth manifold $\mathcal{C}_{p_{1},p_{2},r}^{+}$
Using the ingredients developed in Section 4.1, together with analogies to ancillary results in Section 3, we prove that $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a compact smooth submanifold embedded in $\mathcal{S}_{p,r}^{+}$. To this end, recall the notations in (6). Following the discussion and results in Section 4.1, we shall rewrite $\mathcal{H}_{p_{1},p_{2},r}:=\mathcal{H}_{p_{1},p_{2},r}\setminus\mathcal{V}_{p_{1},p_{2},r}$, and the rest of the notations in (6) are built upon this $\mathcal{H}_{p_{1},p_{2},r}$. By Proposition 7 and a version of Lemma 3.1, $\mathcal{H}_{p_{1},p_{2},r}$ is open in $(\real^{p_{1}\times p_{2}})_{*}^{r}$ and thus has the tangent space $(\real^{p_{1}\times p_{2}})^{r}$. Also, as an analogy to (13), we define the following matrix-valued linear operator $J$ on $\mathcal{H}_{p_{1},p_{2},r}$ by
(22) $J(A):=\left[\begin{array}[]{c}J_{1}(A)\\ J_{2}(A)\\ J_{3}(A)\end{array}\right]=\left[\begin{array}[]{c}\frac{1}{p}(I_{p_{1}^{2}}+K_{(p_{1},p_{1})})[A_{r}\otimes I_{p_{1}},\ldots,A_{r}\otimes I_{p_{1}}]-\frac{2}{p_{1}^{2}p_{2}}\text{vec}(I_{p_{1}})a^{\top}\\ \frac{1}{p}(I_{p_{2}^{2}}+K_{(p_{2},p_{2})})[I_{p_{2}}\otimes A_{r}^{\top},\ldots,I_{p_{2}}\otimes A_{r}^{\top}]-\frac{2}{p_{1}p_{2}^{2}}\text{vec}(I_{p_{2}})a^{\top}\\ 2a^{\top}\end{array}\right],$
where $a=[\text{vec}(A_{1})^{\top},\ldots,\text{vec}(A_{r})^{\top}]^{\top}$. The proof strategy to establish the main result of this section is similar to that in Section 3. The ancillary results to prove Theorem 5 can be established similarly when $r<p$. A slight difference lies in establishing the analogy of Proposition 4; namely, that $\mathcal{D}_{p_{1},p_{2},r}$ is a smooth, closed, and embedded submanifold of $(\real^{p_{1}\times p_{2}})_{*}^{r}$. Following the analogy to Proposition 4, we show that $\operatorname{\text{rank}}(J(A))=\dim\mathcal{R}_{p_{1},p_{2}}$ for any $A\in\mathcal{D}_{p_{1},p_{2},r}$ so that the constant-rank level set theorem applies. As in the proof of Proposition 4, it suffices to verify (14), with $J_{i}:=J_{i}(A)$ for $i=1,2,3$. Now the difference arises in the way showing that $\dim C(J_{1}^{\top})\cap C(J_{2}^{\top})=0$, which relies on the connectivity between row and column factors by canonically indecomopsable matrices in Proposition 8.
@@
Proof 4.8 (Sketch of Proof).
Provided that the first item is true, the last two items directly follow from the argument for the proof of Theorem 5. The results of Lemma 3.1, 3.4, and Proposition 4 can be developed similarly. For Lemma 3.4 with $\mathcal{C}_{p_{1},p_{2},r}$ and $\mathcal{O}_{r}$ instead of $\mathcal{C}_{p_{1},p_{2}}$ and $\mathcal{O}_{p}$, respectively, we additionally show that $\tilde{X}=\varphi_{p_{1},p_{2},r}^{-1}(\varphi_{p_{1},p_{2},r}(X)O)$ is canonically indecomposable for any $X=(X_{1},\ldots,X_{r})\in\mathcal{D}_{p_{1},p_{2},r}$ and $O\in\mathcal{O}_{r}$ so that the action $(O,B)\in\mathcal{O}_{r}\times\mathcal{C}_{p_{1},p_{2},r}\rightarrow BO\in\mathcal{C}_{p_{1},p_{2},r}$ is well-defined.
To prove the first item, recall the operator $J$ in (22). Following the proof of Proposition 4, the first item can be concluded if $C(J_{1}(A)^{\top})\cap C(J_{2}(A)^{\top})=\left\{\mathbf{0}_{pr}\right\}$ for any $A\in\mathcal{D}_{p_{1},p_{2},r}$. This can be done similarly to the proof of Proposition 4, along with the result of Proposition 8.
The tangent spaces of $\mathcal{C}_{p_{1},p_{2},r}$ and $\mathcal{C}_{p_{1},p_{2},r}^{+}$ follow from the proof of Theorem 9.
Proposition 10.
@@
Proof 4.9.
See Apendix A.4 for the proof.
@@
In this section, we prove that $\mathcal{S}_{p}^{++}$ is diffeomorphic to the product manifold $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$ via the map $f:\Sigma\in\mathcal{S}_{p}^{++}\rightarrow(k(\Sigma),c(\Sigma))\in\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$ with its inverse $g:(K,C)\in\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}\rightarrow h(K)Ch(K)^{\top}\in\mathcal{S}_{p}^{++}$. Consequently, we provide a new insight into the smooth structure of $\mathcal{S}_{p}^{++}$ in terms of the separability. This generalizes the result of Proposition $5$ from [hoff2023] on the homeomorphic relationship between $\mathcal{S}_{p}^{++}$ and $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$. We also calculate the differentials of $f$ and $g$ to examine how the tangent vectors transform via the maps $f$ and $g$.
@@
Proof 5.2.
See Appendix A.5 for the proof.
@@
Proof 5.3.
See Appendix A.5 for the proof.
It remains to show that the maps $k$ and $c$ are smooth, proving that the map $f$ is smooth so that the diffeomorphic relationship holds, and compute their differentials. In some sense, the ambiguity due to a constant factor in identifying the factors of the elements in $\mathcal{S}_{p_{1},p_{2}}^{++}$ makes the proof complicated, i.e, $\Sigma_{2}\otimes\Sigma_{1}=(c\Sigma_{2})\otimes(\Sigma_{1}/c)$ for any $c>0$. To avoid this ambiguity, we introduce the orthogonal parameterization of $\mathcal{S}_{p_{1},p_{2}}^{++}$ under $g^{\text{AI}}$ [mccormack2025, simonis2025, cox1987]. Specifically, suppose $\mathcal{E}:=\mathbb{P}(\mathcal{S}_{p_{1}}^{++})\times\mathcal{S}_{p_{2}}^{++}$. Define a diffeomorphism
@@
where $|\Sigma_{1}|=1$. As studied by [mccormack2025, simonis2025], if $\mathcal{S}_{p_{1},p_{2}}^{++}$ is equipped with $g^{\text{AI}}$, the induced metric $\tilde{g}_{\Sigma_{2}\otimes\Sigma_{1}}^{\text{AI}}=\tilde{g}_{1}^{\text{AI}}\oplus\tilde{g}_{2}^{\text{AI}}$ on $\mathcal{E}$ by $\psi_{p_{1},p_{2}}$ via pullback geometry is given by $\tilde{g}_{i}^{\text{AI}}=g_{\Sigma_{i}}^{\text{AI}}/p_{-i}$ for $p_{-1}=p_{2}$ and $p_{-2}=p_{1}$. Then for the Kronecker map $k$, let $\eta_{p_{1},p_{2}}:=\psi_{p_{1},p_{2}}\circ k$. Since $k:=\psi_{p_{1},p_{2}}^{-1}\circ\eta_{p_{1},p_{2}}$, if $\eta_{p_{1},p_{2}}$ is smooth, so is $k$. Also, the chain rule implies that
(23) $dk(\Sigma)[V]=d\psi_{p_{1},p_{2}}^{-1}(k(\Sigma))[d\eta_{p_{1},p_{2}}(\Sigma)[V]]$
@@
(24) $d\psi_{p_{1},p_{2}}^{-1}(\Sigma_{1},\Sigma_{2}):(U_{1},U_{2})\in T_{(\Sigma_{1},\Sigma_{2})}\mathcal{E}\rightarrow U_{2}\otimes\Sigma_{1}+\Sigma_{2}\otimes U_{1}\in T_{\Sigma_{2}\otimes\Sigma_{1}}\mathcal{S}_{p_{1},p_{2}}^{++}.$
After identifying $d\eta_{p_{1},p_{2}}(\Sigma)[V]$ via the manifold implicit function theorem ([loomis2014], Section $3.11$), we use (23)–(24) to obtain $dk(\Sigma)[V]$ and so $dc(\Sigma)[V]$ using Lemma 5.1.
Proposition 12.
The Kronecker map $k:\mathcal{S}_{p}^{++}\rightarrow\mathcal{S}_{p_{1},p_{2}}^{++}$ is smooth. Consequently, the map $f:\Sigma\in\mathcal{S}_{p}^{++}\rightarrow(k(\Sigma),c(\Sigma))\in\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$ is so for either $h\in\mathcal{S}_{p_{1},p_{2}}^{++}$ or $h\in\mathcal{C}_{p_{1},p_{2}}^{++}$. Therefore, $\mathcal{S}_{p}^{++}$ is diffeomorphic to $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$ as the map $g$ in Proposition 11 is also smooth. Moreover, let $k(\Sigma)=\Sigma_{2}\otimes\Sigma_{1}$ with $|\Sigma_{1}|=1$, $c(\Sigma)=C$, and $V\in T_{\Sigma}\mathcal{S}_{p}^{++}\equiv\mathcal{S}_{p}$. Also, define the bilinear operator $\mathcal{R}_{C}:T_{(\Sigma_{1},\Sigma_{2})}\mathcal{E}\rightarrow T_{(\Sigma_{1},\Sigma_{2})}\mathcal{E}$ by
@@
Proof 5.4.
See Appendix A.5 for the proof.

### v3

@@ 4.1 Canonically decomposable matrices
As an example of canonical decomposability, for generic element $(A_{1},A_{2})$ in $(\real^{4\times 7})^{2}$, there exists a $(P,Q)\in GL_{4}\times GL_{7}$ such that $PA_{i}Q^{-1}=\oplus_{j=1}^{3}B_{ij}$, where $B_{i1},B_{i2}\in\real^{1\times 2}$ and $B_{i3}\in\real^{2\times 3}$ ([19], Example $2.8$). Here the term generic should be understood as an almost sure sense. An example with specific values of $A_{i}$’s and $(P,Q)$ is provided in Example $4$ of [43].
@@
The notion of canonical decomposability is mainly motivated by Kronecker quiver representation and its applications in the analysis of the sample size threshold for the existence of Kronecker MLE [19] (see [36, 37] also). To say informally, the $n$-Kronecker quiver $Q$ is a directed acyclic graph of two vertices $x$ and $y$ with $n$ arrows. Then, a representation of $Q$ is to assign a finite-dimensional vector space to each vertex. If this representation cannot be written as a direct sum of a non-trivial subrepresentation, such a representation is referred to as a $\sigma$-stable representation. In the context of the Kronecker MLE problem, these vector spaces correspond to $\real{}^{p_{2}}$ and $\real{}^{p_{1}}$, and the arrows correspond to the $n$ data matrices. Using this Kronecker-quiver representation, along with the group-invariant theory, [19] characterized the scenarios of $(p_{1},p_{2},r)$ for which Kronecker MLE exists (see their Theorem $1.2$). For generic $(p_{1},p_{2})$, it turns out that $r>p_{1}/p_{2}+p_{2}/p_{1}$. Under this threshold, the uniqueness of Kronecker MLE also follows if it exists.
@@
To interpret this threshold, the canonical decomposability of $r$ data matrices in Definition 4.1 corresponds to whether the $r$ data matrices induce a $\sigma$-stable Kronecker quiver representation as arrows (see Section $3$ and $5$ of [19]). Then the threshold on $r$ for which Kronecker MLE exists comes from the minimum number of arrows for which the factors in $\real{}^{p_{1}}$ and $\real{}^{p_{2}}$ are well-connected. We shall formally formulate this below. We emphasize that this decomposability notion also appears in other works on the sample size threshold analysis for Kronecker MLE. For example, [61] studied the threshold by analyzing the contribution of each block in the canonical decomposition of data matrices to the growth of the objective function $d$ in (3) (see Section $6.6$ of [61]). They referred canonically decomposable data matrices to as bad samples.
@@
In these works, the sample size threshold should be understood in a generic (almost sure) sense. If $r\geq p$ and $r$ data matrices are linearly independent, then the Kronecker MLE always uniquely exists [62], not just generically. However, in a strict algebra sense, even if $r>p_{1}/p_{2}+p_{2}/p_{1}$ and the linear independence holds for data matrices, the Kronecker MLE may not exist as observed in Example 2.1. Nevertheless, note that the data matrices in that example are canonically decomposable. Thus, a natural question one could raise is whether any canonically decomposable matrices never admit the Kronecker MLE. It turns out that the answer is no, as shown in Example 4.1. This example also suggests that the canonically decomposable matrices are the singularities that may prevent the set of rank-$r$ cores from being a smooth manifold, in light of Sard’s theorem.
@@
Denote the subset of canonically decomposable matrices in $(\real^{p_{1}\times p_{2}})^{r}$ by $\mathcal{V}_{p_{1},p_{2},r}$. We show that this set is closed and has a Lebesgue measure zero. Note that its analogous results have been proven based on group-invariant theory and representation-theoretic approaches (see Proposition $3.19$ of [34], and Lemma $2.16$ and Section $5$ of [19]). However, we provide a proof using a more direct language of algebraic geometry to make the article self-contained and better motivate the canonical decomposability in studying the fixed-rank core covariance manifold.
@@
Proof.
See Appendix A.3 for the proof. ∎
Proposition 4.3.
@@
Proof.
See Appendix A.3 for the proof. ∎
@@
Proof.
Suppose otherwise. Then there exists indecomopsable $A$ and $(P,Q)\in GL_{p_{1}}\times GL_{p_{2}}$ such that the graph $G_{A,P,Q}$ is disconnected. Hence, there exist partitions $U_{1}$ and $U_{2}$ of $\left\{s_{i}\right\}$ and accordingly $W_{1}$ and $W_{2}$ of $\left\{q_{j}\right\}$ such that a vertex in $U_{1}$ (resp. $U_{2}$) is never connected to $W_{2}$ (resp. $W_{1}$). After arranging the row and columns of $PA_{i}Q^{-1}$, we can obtain $(P^{\prime},Q^{\prime})\in GL_{p_{1}}\times GL_{p_{2}}$ such that $P^{\prime}A_{i}(Q^{\prime})^{-1}$ is of non-trivial block-diagonal form where the zero entries correspond to the absence of edges between $U_{1}$ (resp. $U_{2}$) and $W_{2}$ (resp. $W_{1}$), contradicting the indecomposability of $A$. ∎
@@ 4.2 Proof of the smooth manifold $\mathcal{C}_{p_{1},p_{2},r}^{+}$
Using the ingredients developed in Section 4.1, together with analogies to ancillary results in Section 3, we prove that $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a compact smooth submanifold embedded in $\mathcal{S}_{p,r}^{+}$. To this end, recall the notations in (7). Following the discussion and results in Section 4.1, we shall rewrite $\mathcal{H}_{p_{1},p_{2},r}:=\mathcal{H}_{p_{1},p_{2},r}\setminus\mathcal{V}_{p_{1},p_{2},r}$, and the rest of the notations in (7) are built upon this $\mathcal{H}_{p_{1},p_{2},r}$. By Proposition 4.3 and a version of Lemma 3.1, $\mathcal{H}_{p_{1},p_{2},r}$ is open in $(\real^{p_{1}\times p_{2}})_{*}^{r}$ and thus has the tangent space $(\real^{p_{1}\times p_{2}})^{r}$. Also, as an analogy to (14), we define the following matrix-valued linear operator $J$ on $\mathcal{H}_{p_{1},p_{2},r}$ by
$J(A):=\left[\begin{array}[]{c}J_{1}(A)\\ J_{2}(A)\\ J_{3}(A)\end{array}\right]=\left[\begin{array}[]{c}\frac{1}{p}(I_{p_{1}^{2}}+K_{(p_{1},p_{1})})[A_{r}\otimes I_{p_{1}},\ldots,A_{r}\otimes I_{p_{1}}]-\frac{2}{p_{1}^{2}p_{2}}\text{vec}(I_{p_{1}})a^{\top}\\ \frac{1}{p}(I_{p_{2}^{2}}+K_{(p_{2},p_{2})})[I_{p_{2}}\otimes A_{r}^{\top},\ldots,I_{p_{2}}\otimes A_{r}^{\top}]-\frac{2}{p_{1}p_{2}^{2}}\text{vec}(I_{p_{2}})a^{\top}\\ 2a^{\top}\end{array}\right],$ (22)
where $a=[\text{vec}(A_{1})^{\top},\ldots,\text{vec}(A_{r})^{\top}]^{\top}$. The proof strategy to establish the main result of this section is similar to that in Section 3. The ancillary results to prove Theorem 3.5 can be established similarly when $r<p$. A slight difference lies in establishing the analogy of Proposition 3.2; namely, that $\mathcal{D}_{p_{1},p_{2},r}$ is a smooth, closed, and embedded submanifold of $(\real^{p_{1}\times p_{2}})_{*}^{r}$. Following the analogy to Proposition 3.2, we show that $\operatorname{\text{rank}}(J(A))=\dim\mathcal{R}_{p_{1},p_{2}}$ for any $A\in\mathcal{D}_{p_{1},p_{2},r}$ so that the constant-rank level set theorem applies. As in the proof of Proposition 3.2, it suffices to verify (15), with $J_{i}:=J_{i}(A)$ for $i=1,2,3$. Now the difference arises in the way showing that $\dim C(J_{1}^{\top})\cap C(J_{2}^{\top})=0$, which relies on the connectivity between row and column factors by canonically indecomopsable matrices in Proposition 4.5.
@@
Sketch of Proof.
Provided that the first item is true, the last two items directly follow from the argument for the proof of Theorem 3.5. The results of Lemma 3.1, 3.3, and Proposition 3.2 can be developed similarly. For Lemma 3.3 with $\mathcal{C}_{p_{1},p_{2},r}$ and $\mathcal{O}_{r}$ instead of $\mathcal{C}_{p_{1},p_{2}}$ and $\mathcal{O}_{p}$, respectively, we additionally show that $\tilde{X}=\varphi_{p_{1},p_{2},r}^{-1}(\varphi_{p_{1},p_{2},r}(X)O)$ is canonically indecomposable for any $X=(X_{1},\ldots,X_{r})\in\mathcal{D}_{p_{1},p_{2},r}$ and $O\in\mathcal{O}_{r}$ so that the action $(O,B)\in\mathcal{O}_{r}\times\mathcal{C}_{p_{1},p_{2},r}\rightarrow BO\in\mathcal{C}_{p_{1},p_{2},r}$ is well-defined.
To prove the first item, recall the operator $J$ in (22). Following the proof of Proposition 3.2, the first item can be concluded if $C(J_{1}(A)^{\top})\cap C(J_{2}(A)^{\top})=\left\{\mathbf{0}_{pr}\right\}$ for any $A\in\mathcal{D}_{p_{1},p_{2},r}$. This can be done similarly to the proof of Proposition 3.2, along with the result of Proposition 4.5. ∎
The tangent spaces of $\mathcal{C}_{p_{1},p_{2},r}$ and $\mathcal{C}_{p_{1},p_{2},r}^{+}$ follow from the proof of Theorem 4.6.
Proposition 4.7.
@@
Proof.
See Apendix A.4 for the proof. ∎
@@
In this section, we prove that $\mathcal{S}_{p}^{++}$ is diffeomorphic to the product manifold $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$ via the map $f:\Sigma\in\mathcal{S}_{p}^{++}\rightarrow(k(\Sigma),c(\Sigma))\in\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$ with its inverse $g:(K,C)\in\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}\rightarrow h(K)Ch(K)^{\top}\in\mathcal{S}_{p}^{++}$. Consequently, we provide a new insight into the smooth structure of $\mathcal{S}_{p}^{++}$ in terms of the separability. This generalizes the result of Proposition $5$ from [33] on the homeomorphic relationship between $\mathcal{S}_{p}^{++}$ and $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$. We also calculate the differentials of $f$ and $g$ to examine how the tangent vectors transform via the maps $f$ and $g$.
@@
Proof.
See Appendix A.5 for the proof. ∎
@@
Proof.
See Appendix A.5 for the proof. ∎
It remains to show that the maps $k$ and $c$ are smooth, proving that the map $f$ is smooth so that the diffeomorphic relationship holds, and compute their differentials. In some sense, the ambiguity due to a constant factor in identifying the factors of the elements in $\mathcal{S}_{p_{1},p_{2}}^{++}$ makes the proof complicated, i.e, $\Sigma_{2}\otimes\Sigma_{1}=(c\Sigma_{2})\otimes(\Sigma_{1}/c)$ for any $c>0$. To avoid this ambiguity, we introduce the orthogonal parameterization of $\mathcal{S}_{p_{1},p_{2}}^{++}$ under $g^{\text{AI}}$ [46, 59, 17]. Specifically, suppose $\mathcal{E}:=\mathbb{P}(\mathcal{S}_{p_{1}}^{++})\times\mathcal{S}_{p_{2}}^{++}$. Define a diffeomorphism
@@
where $|\Sigma_{1}|=1$. As studied by [46, 59], if $\mathcal{S}_{p_{1},p_{2}}^{++}$ is equipped with $g^{\text{AI}}$, the induced metric $\tilde{g}_{\Sigma_{2}\otimes\Sigma_{1}}^{\text{AI}}=\tilde{g}_{1}^{\text{AI}}\oplus\tilde{g}_{2}^{\text{AI}}$ on $\mathcal{E}$ by $\psi_{p_{1},p_{2}}$ via pullback geometry is given by $\tilde{g}_{i}^{\text{AI}}=g_{\Sigma_{i}}^{\text{AI}}/p_{-i}$ for $p_{-1}=p_{2}$ and $p_{-2}=p_{1}$. Then for the Kronecker map $k$, let $\eta_{p_{1},p_{2}}:=\psi_{p_{1},p_{2}}\circ k$. Since $k:=\psi_{p_{1},p_{2}}^{-1}\circ\eta_{p_{1},p_{2}}$, if $\eta_{p_{1},p_{2}}$ is smooth, so is $k$. Also, the chain rule implies that
$dk(\Sigma)[V]=d\psi_{p_{1},p_{2}}^{-1}(k(\Sigma))[d\eta_{p_{1},p_{2}}(\Sigma)[V]]$ (23)
@@
$d\psi_{p_{1},p_{2}}^{-1}(\Sigma_{1},\Sigma_{2}):(U_{1},U_{2})\in T_{(\Sigma_{1},\Sigma_{2})}\mathcal{E}\rightarrow U_{2}\otimes\Sigma_{1}+\Sigma_{2}\otimes U_{1}\in T_{\Sigma_{2}\otimes\Sigma_{1}}\mathcal{S}_{p_{1},p_{2}}^{++}.$ (24)
After identifying $d\eta_{p_{1},p_{2}}(\Sigma)[V]$ via the manifold implicit function theorem ([41], Section $3.11$), we use (23)–(24) to obtain $dk(\Sigma)[V]$ and so $dc(\Sigma)[V]$ using Lemma 5.1.
Proposition 5.3.
The Kronecker map $k:\mathcal{S}_{p}^{++}\rightarrow\mathcal{S}_{p_{1},p_{2}}^{++}$ is smooth. Consequently, the map $f:\Sigma\in\mathcal{S}_{p}^{++}\rightarrow(k(\Sigma),c(\Sigma))\in\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$ is so for either $h\in\mathcal{S}_{p_{1},p_{2}}^{++}$ or $h\in\mathcal{C}_{p_{1},p_{2}}^{++}$. Therefore, $\mathcal{S}_{p}^{++}$ is diffeomorphic to $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$ as the map $g$ in Proposition 5.2 is also smooth. Moreover, let $k(\Sigma)=\Sigma_{2}\otimes\Sigma_{1}$ with $|\Sigma_{1}|=1$, $c(\Sigma)=C$, and $V\in T_{\Sigma}\mathcal{S}_{p}^{++}\equiv\mathcal{S}_{p}$. Also, define the bilinear operator $\mathcal{R}_{C}:T_{(\Sigma_{1},\Sigma_{2})}\mathcal{E}\rightarrow T_{(\Sigma_{1},\Sigma_{2})}\mathcal{E}$ by
@@
Proof.
See Appendix A.5 for the proof. ∎